NOAA GOES SATOPS Morning Report August 21, 2003 Day 233

Status Report From: National Oceanic and Atmospheric Administration
Posted: Thursday, August 21, 2003


The IPACS disk replacement effort that took place yesterday occurred without event.  PEAK to extend their control operations 4 extra passes again today to accommodate the second phase of IPACS disk replacement work at SOCC.

All DMSP operations were nominal over the past 24 hours.


All POES operations were nominal over the past 24 hours.


At 1333z on August 20, a GOES-9 (sounding) frame break was observed and attributed to high BER (unknown).

At 1735z on August 20 and 0254z on August 21, GOES-12 frame breaks were observed and attributed to loss of signal synchronization.

GOES battery #2 reconditioning is in process for all five spacecraft as part of normal pre-eclipse season preparations.  For GOES Fall eclipse and KOZ schedules, please reference: .

GOES-8 continues its westward drift at 1.06 deg/day. The drift table with the estimated current position can be found at  GOES-8 will be stopped at 165E on August 24 at 0930z.

All GOES operations were nominal over the past 24 hours.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter